OpenStudy (anonymous):

Optimization problem: A sector of angle θ is cut from a circle of radius 9 inches and the resulting edges are brought together to form a cone. Find the magnitude of θ so that volume of the cone is maximized.

OpenStudy (anonymous):

If you're also seeing a line of question marks... it's supposed to be \(\theta\).

OpenStudy (anonymous):

Hint: you have two equivalent expressions for area (\(A\)) in terms of two different variables; one which you find in the volume equation (\(r\)), and the other is the variable you're optimizing (\(\theta\)). You also have an expression that relates the other unknown in the volume eq (\(h\)) in terms of the radius (\(r\)), which in turn can be expressed in terms of the angle (\(\theta\)).

OpenStudy (anonymous):

All this means you can write volume in terms of the angle.

OpenStudy (anonymous):

yeah, Thanks! Had a hard time writing the volume in terms of angle.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Oh hold up, I made a mistake somewhere along the line! The \(r\) you see above the circle is the arc length, but is not the same as the \(r\) on the cone (\(r\) is NOT the cone's radius). Let me draw that again with correct variables: